# The action of an inverse semigroup on its Stone-Čech compactification

### v1

@@ 3. The Stone-Čech transformation groupoid of an inverse semigroup
Sets of this form generate the topology on $\beta\mathcal{S}$. Each element $s\in\mathcal{S}$ determines a principal ultrafilter $\{A\subseteq\mathcal{S}:s\in A\}=\{s\}^{\subseteq}$; so the canonical embedding described in Example 2.2 is given by $\iota(s)=\{s\}^{\subseteq}$. In what follows we will drop the $\iota$ and consider $\mathcal{S}$ as a subset of $\beta\mathcal{S}$. We then have that $\overline{\mathcal{S}}=\beta\mathcal{S}$. Furthermore, it is a straightforward exercise to show that
$U_{A}=\overline{A},$
that is, the basic open set $U_{A}$ is the closure of $A$ in $\beta\mathcal{S}$.
Since $\mathcal{S}$ is a discrete space, there is a canonical homeomorphism $f:\overline{A}\to\beta A$. With ultrafilters, this map is defined by $f(\xi)=\{B\in\xi:B\subseteq A\}$.
Theorem 3.1.
The reduced groupoid C-algebra $\textup{C}^{*}_{r}(\mathcal{S}\ltimes\beta\mathcal{S})$ is isomorphic to the reduced crossed product $\ell^{\infty}(\mathcal{S})\rtimes_{r}\mathcal{S}$. In particular, $\mathcal{S}\ltimes\beta\mathcal{S}$ is amenable if and only if $\textup{C}^{*}_{r}(\mathcal{S})$ is exact.
Proof.
[…]
In what remains of this paper we study the action of $\mathcal{S}$ on its Stone-Čech compactification and properties of its transformation groupoid $\mathcal{S}\ltimes\beta\mathcal{S}$. In [EP16], Exel and Pardo proved many results about such actions and groupoids, so we recall some terminology and facts from there.
@@
Elements of $F_{s}$ are the fixed points of $s$ and the elements of $TF_{s}$ are the trivially fixed points of $s$; this is the terminology given in [EP16].
@@
Let $s\in\mathcal{S}$. If $\lambda_{s}\xi=\xi$ for some $\xi\in\overline{s^{*}\mathcal{S}}$, then there is a subset $A\subseteq s^{*}\mathcal{S}$ such that $\lambda_{s}\eta=\eta$, for all $\eta\in\overline{A}$. In particular, $\textup{Iso}(\mathcal{S}\ltimes\beta\mathcal{S})$ is open.
@@
The result is plain if $\xi$ is a principal ultrafilter, so we may assume that $\xi$ is a free ultrafilter (and therefore that $s^{*}\mathcal{S}$ is infinite.) Let $B\subset s^{*}\mathcal{S}$ be an infinite subset with infinite complement. Either $B\in\xi$ or $s^{*}\mathcal{S}\backslash B\in\xi$; without loss of generality, take $B\in\xi$. Because $\lambda_{s}:s^{*}\mathcal{S}\to s\mathcal{S}$ is a bijection, $sB\subset s\mathcal{S}$ is also an infinite subset with infinite complement. Because $\B\mathcal{S}\backslash$ B and $\sB\mathcal{S}\backslash$ sB are both countably infinite, we may extend the map $\lambda_{s}:B\to s^{*}B$ to a homeomorphism $f:\beta\mathcal{S}\to\beta\mathcal{S}$ with the property that $f(\xi)=\xi$. By Frolík’s theorem ([Frolik1971], Theorem 3.1), there is a basic open neighbourhood $U=\overline{C}$ of $\xi$ ($C\subseteq\mathcal{S}$) such that $f(\eta)=\eta$ for all $\eta\in U$. The result follows after taking $A=B\cap C$. ∎
@@
Suppose that there is some $\xi\in\overline{s^{*}\mathcal{S}}$ such that $\lambda_{s}\xi=\xi$. By Lemma 3.3, there exists $A\subseteq s^{*}\mathcal{S}$ such that $U_{A}=\overline{A}$ is fixed by $\lambda_{s}$. In particular, we have $sa=a$, for all $a\in A$, so $saa^{*}=aa^{*}$ for all $a\in A$. Said differently, $aa^{*}\in\mathcal{J}_{s}$, for all $a\in A$. Thus,
$A\subseteq\bigcup_{a\in A}a\mathcal{S}=\bigcup_{a\in A}aa^{*}\mathcal{S}\subseteq\bigcup_{e\in\mathcal{J}}e\mathcal{S}\implies\bigcup_{e\in\mathcal{J}}e\mathcal{S}\in\xi,$
because $\xi$ is a filter and filters are upwards closed. Thus $\xi\in\overline{\bigcup_{e\in\mathcal{J}}e\mathcal{S}}$, and so $F_{s}\subseteq\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}$. Reverse containment follows from the continuity of $\lambda_{s}$, completing the proof. ∎
@@
Let $\mathcal{G}$ be a groupoid. We say that
@@
    $\mathcal{G}$ is principal if $g\in\mathcal{G}$ and $s(g)=r(g)$ implies $g\in\mathcal{G}^{(0)}$,
@@
We have that $sx=x$ and so $sxx^{*}=xx^{*}$. Since $x\in xS=xx^{*}\mathcal{S}$, we have that $xx^{*}\mathcal{S}\in x$. Thus $s,x]=[sxx^{*},x]=[xx^{*},x]\in(\mathcal{S}\ltimes\beta\mathcal{S})^{(0)}$. Because principal filters are dense in $(\mathcal{S}\ltimes\beta\mathcal{S})^{(0)}$, $\mathcal{S}\ltimes\beta\mathcal{S}$ is essentially principal. ∎
@@
$(1)\implies(2)$ By [EP16], $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff if and only if $TF_{s}=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}$ is closed relative to $\overline{s^{*}\mathcal{S}}$. Lemma 3.4 says that $F_{s}$ is the smallest closed set that contains $e\mathcal{S}$ for all $e\in\mathcal{J}_{s}$, so $F_{s}\subseteq TF_{s}$, and hence $F_{s}=TF_{s}$. Because every fixed point is trivially fixed, it follows from Lemma 3.2 that $\mathcal{S}\ltimes\beta\mathcal{S}$ is principal.
@@
$(1)\implies(2)$ By Theorem 3.7, if $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff, then $TF_{s}=F_{s}=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}$ for all $s\in\mathcal{S}$. This set is compact and each $\overline{e\mathcal{S}}$ is open, so there is a finite $F\subseteq\mathcal{J}_{s}$ such that
@@
where the last equality follows as $F$ is finite. Because elements of $\mathcal{S}$ are isolated in $\beta\mathcal{S}$, this implies $\bigcup_{f\in F}f\mathcal{S}=\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}$.
@@
implying $F_{s}=TF_{s}$. By Theorem 3.7, $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff. ∎
@@
An inverse semigroup is an abstract pseudogroup if it is a complete, infinitely distributive inverse monoid (see [lawson]). The symmetric inverse monoid $\mathcal{I}(X)$ is an abstract pseudogroup. An inverse semigroup with zero $\mathcal{S}$ is $E^{*}$-unitary if $s\in\mathcal{S}$ and $\mathcal{J}_{s}\neq\{0\}$ implies $s\in E_{S}$. Graph inverse semigroups and free inverse semigroups are $E^{*}$-unitary [L]. The next result is an immediate consequence of Theorem 3.8.
@@
Every transformation groupoid of an $E^{*}$-unitary inverse semigroup is Hausdorff by [EP16]. Alternatively, observe that if $\mathcal{J}_{s}$ is nonzero, then $s\in\mathcal{J}_{s}$, and thus $\mathcal{J}_{s}=\{s\}^{\geqslant}$. ∎
The main result of [starling] says that the action of $\mathcal{S}$ on its Exel spectrum is amenable if and only if every action of $\mathcal{S}$ is amenable. Our last result has a similar flavour, and provides some clear obstructions to the Hausdorffness of $\mathcal{S}\ltimes\beta\mathcal{S}$.
@@
By assumption, each $D^{\alpha}_{f}$ is closed. Since $F$ is finite, $TF^{\alpha}_{s}$ is closed, so by [EP16, Theorem 3.15], $\mathcal{S}\ltimes\mathcal{X}$ is Hausdorff. ∎

### v2

@@ 3. The Stone-Čech transformation groupoid of an inverse semigroup
Sets of this form generate the topology on $\beta\mathcal{S}$. Each element $s\in\mathcal{S}$ determines a principal ultrafilter $\{A\subseteq\mathcal{S}:s\in A\}=\{s\}^{\subseteq}$, so the canonical embedding described in Example 2.1 is given by $\iota(s)=\{s\}^{\subseteq}$. In what follows we will drop the $\iota$ and consider $\mathcal{S}$ as a subset of $\beta\mathcal{S}$. We then have that $\overline{\mathcal{S}}=\beta\mathcal{S}$. Now observe that
$U_{A}=\overline{A};$
that is, the basic open set $U_{A}$ is the closure of $A$ in $\beta\mathcal{S}$. To see this, note that $\xi\in\overline{A}$ if and only if, for all $B\in\xi$, there is some $a\in A$ with $a\in B$. This condition forces $A\in\xi$, because $\xi$ is an ultrafilter and otherwise one would have $\mathcal{S}\backslash A\in\xi$. Conversely, if $A\in\xi$, then already $A\cap B\neq\emptyset$ for all $B\in\xi$, so $\xi\in\overline{A}$.
Since $\mathcal{S}$ is a discrete space, there is a canonical homeomorphism $f:\beta\mathcal{S}\supset\overline{A}\to\beta A$, for all $A\subset\mathcal{S}$. With ultrafilters, this map is defined by $f(\xi)=\{B\in\xi:B\subseteq A\}$.
We now turn to studying the action of $\mathcal{S}$ on its Stone-Čech compactification and properties of its transformation groupoid $\mathcal{S}\ltimes\beta\mathcal{S}$. In [7], Exel and Pardo proved many general results about such actions and groupoids, so we recall some terminology and facts from there.
@@
Following the terminology given in [7], the elements of $F_{s}$ are called the fixed points of $s$, and the elements of $TF_{s}$ are called the trivially fixed points of $s$.
@@
Consequently, $\mathcal{S}\ltimes\beta\mathcal{S}$ is principal if and only if $TF_{s}=F_{s}$ for all $s\in\mathcal{S}$.
@@
Let $s\in\mathcal{S}$. If $\lambda_{s}\xi=\xi$ for some $\xi\in\overline{s^{*}\mathcal{S}}$, then there is a nonempty subset $A\subseteq s^{*}\mathcal{S}$ such that $A\in\xi$ and $\lambda_{s}\eta=\eta$, for all $\eta\in\overline{A}$. In particular, $\textup{Iso}(\mathcal{S}\ltimes\beta\mathcal{S})$ is nonempty and open.
@@
The result is plain if $\xi$ is a principal ultrafilter, so we may assume that $\xi$ is a free ultrafilter (and therefore that $s^{*}\mathcal{S}$ is infinite.) Let $B\subset s^{*}\mathcal{S}$ be an infinite subset with infinite complement. Either $B\in\xi$ or $s^{*}\mathcal{S}\backslash B\in\xi$; without loss of generality, take $B\in\xi$. Because $\lambda_{s}:s^{*}\mathcal{S}\to s\mathcal{S}$ is a bijection, $sB\subset s\mathcal{S}$ is also an infinite subset with infinite complement. Because $\B\mathcal{S}\backslash$ B and $\sB\mathcal{S}\backslash$ sB are both countably infinite, we may extend the map $\lambda_{s}:B\to sB$ to a homeomorphism $f:\beta\mathcal{S}\to\beta\mathcal{S}$ with the property that $f(\xi)=\xi$. By Frolík’s theorem ([10], Theorem 3.1), there is a basic open neighbourhood $U=\overline{C}$ of $\xi$ ($C\subseteq\mathcal{S}$) such that $f(\eta)=\eta$ for all $\eta\in U$. The result follows after taking $A=B\cap C$. ∎
@@
Suppose that there is some $\xi\in\overline{s^{*}\mathcal{S}}$ such that $\lambda_{s}\xi=\xi$. By Lemma 3.2, there exists $A\subseteq s^{*}\mathcal{S}$ with $A\in\xi$ such that $U_{A}=\overline{A}$ is fixed by $\lambda_{s}$. In particular, we have $sa=a$, for all $a\in A$, so $saa^{*}=aa^{*}$ for all $a\in A$. Said differently, $aa^{*}\in\mathcal{J}_{s}$, for all $a\in A$. Thus,
$A\subseteq\bigcup_{a\in A}a\mathcal{S}=\bigcup_{a\in A}aa^{*}\mathcal{S}\subseteq\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}\implies\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}\in\xi,$
because $\xi$ is a filter and filters are upwards closed. Thus $\xi\in\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}$, and so $F_{s}\subseteq\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}$. Reverse containment follows from the continuity of $\lambda_{s}$, completing the proof. ∎
@@
Let $\mathcal{G}$ be a topological groupoid. We say that
@@
    $\mathcal{G}$ is principal if $g\in\mathcal{G}$ and $\mathbf{d}(g)=\mathbf{r}(g)$ implies $g\in\mathcal{G}^{(0)}$,
@@
We point out to the reader that some authors use the term topologically free to mean what we have defined here as essentially principal.
@@
We have that $sx=x$ and so $sxx^{*}=xx^{*}$. Since $x\in x\mathcal{S}=xx^{*}\mathcal{S}$, we have that $xx^{*}\mathcal{S}\in\{x\}^{\subseteq}$. Thus $s,x]=[sxx^{*},x]=[xx^{*},x]\in(\mathcal{S}\ltimes\beta\mathcal{S})^{(0)}$. Because principal filters are dense in $(\mathcal{S}\ltimes\beta\mathcal{S})^{(0)}$, $\mathcal{S}\ltimes\beta\mathcal{S}$ is essentially principal. ∎
@@
$(1)\implies(2)$ By [7], $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff if and only if $TF_{s}=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}$ is closed relative to $\overline{s^{*}\mathcal{S}}$. Because $\overline{s^{*}\mathcal{S}}$ is already closed, this is equivalent to $TF_{s}$ being closed. Lemma 3.3 says that $F_{s}$ is the smallest closed set that contains $e\mathcal{S}$ for all $e\in\mathcal{J}_{s}$, so $F_{s}\subseteq TF_{s}$, and hence $F_{s}=TF_{s}$. Because every fixed point is trivially fixed, it follows from Lemma 3.1 that $\mathcal{S}\ltimes\beta\mathcal{S}$ is principal.
@@
$(1)\implies(2)$ By Theorem 3.6, if $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff, then it is principal and so $F_{s}=TF_{s}$ for all $s\in\mathcal{S}$. We always have $TF_{s}=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}$ and Lemma 3.3 implies $F_{s}=\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}$. Thus for all $s\in\mathcal{S}$ we have $\overline{\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}}=\bigcup_{e\in\mathcal{J}_{s}}\overline{e\mathcal{S}}$. This set is compact and each $\overline{e\mathcal{S}}$ is open, so there is a finite $F\subseteq\mathcal{J}_{s}$ such that
@@
where the first equality is the expression $F_{s}=TF_{s}$ and the last equality follows as $F$ is finite. Because elements of $\mathcal{S}$ are isolated in $\beta\mathcal{S}$, this implies $\bigcup_{f\in F}f\mathcal{S}=\bigcup_{e\in\mathcal{J}_{s}}e\mathcal{S}$.
@@
where the middle equality follows because the union in question is finite. This implies $F_{s}=TF_{s}$. By Theorem 3.6, $\mathcal{S}\ltimes\beta\mathcal{S}$ is Hausdorff. ∎
@@
An inverse semigroup is an abstract pseudogroup if it is a complete, infinitely distributive inverse monoid (see [12]). This notion is not to be confused with other inequivalent definitions of pseudogroup, such as a pseudogroup in the sense of Renault (cf. [20]). The symmetric inverse monoid $\mathcal{I}(X)$ is an abstract pseudogroup. An inverse semigroup with zero $\mathcal{S}$ is $E^{*}$-unitary if $s\in\mathcal{S}$ and $\mathcal{J}_{s}\neq\{0\}$ implies $s\in E_{S}$. Graph inverse semigroups and free inverse semigroups are $E^{*}$-unitary [13]. The next result is an immediate consequence of Theorem 3.7.
@@
Every transformation groupoid of an $E^{*}$-unitary inverse semigroup is Hausdorff by [7]. Alternatively, observe that if $\mathcal{J}_{s}$ is nonzero, then $s\in\mathcal{J}_{s}$, and thus $\mathcal{J}_{s}=\{s\}^{\geqslant}$. ∎
The main result of [8] says that the action of $\mathcal{S}$ on its spectrum $\widehat{E}_{0}$ (the space of nonzero semicharacters on $E_{\mathcal{S}}$) is amenable if and only if every action of $\mathcal{S}$ is amenable. Our last result in this section has a similar flavour, and provides some clear obstructions to the Hausdorffness of $\mathcal{S}\ltimes\beta\mathcal{S}$.
@@
By assumption, each $D^{\alpha}_{f}$ is closed. Since $F$ is finite, $TF^{\alpha}_{s}$ is closed, so by [7]*Theorem 3.15, $\mathcal{S}\ltimes\mathcal{X}$ is Hausdorff. ∎